# Weaving the (AdS) spaces with partial entanglement entropy threads

### v3

@@ 3 Reformulation of the RT formula
$\mathcal{N}(\Sigma_{A})=\frac{1}{2}\int_{\Sigma_{A}}\mathrm{d}\Sigma_{A}\sqrt{h}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}|V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)|\,.$ (23)
Here $\mathrm{d}\Sigma_{A}\sqrt{h}$ is the area of an infinitesimal area element on $\Sigma_{A}$, and the coefficient $1/2$ appears because we integrate x over the whole boundary such that the PEE thread connecting any two boundary point is doubly counted. The equivalence between (23) and (22) is guaranteed by the divergenceless property of the vector fields $V_{\textbf{x}}^{\mu}$.
According to (23), the density of intersections between $\Sigma_{A}$ and the PEE network at point $Q$ is just given by
density of $\textit{density of intersections}~{}~{}=~{}~{}\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}|V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)|\,.$ (24)
Our task then reduces to calculating the above integration and prove
$\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}|V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)|=\frac{1}{4G}\,.$ (25)
Since (25) claims that the integration is independent both of the point $Q$ and the surface $\Sigma_{A}$ in which $Q$ is embedded, proving (25) equals to proving that, the density of intersection numbers for any bulk area element $\mathrm{d}\Sigma$ is $1/(4G)$.
Although the analysis for static spherical regions [1] looks quite special, it contains the key ingredient to prove (25). Let us consider an arbitrary area element $(\mathrm{d}\Sigma$ at any point $Q$ in the bulk and the associated normal unit vector $n^{\mu}(Q)$, directly calculating the integration (24) is complicated, because of the absolute symbol in the integrand. Our strategy is to find the boundary region $\mathcal{B}$ such that, for any $\textbf{x}\in\mathcal{B}$ we have
@@
density of $\textit{density of intersections}~{}~{}=~{}~{}\int_{\mathcal{B}}\mathrm{d}^{d-1}\textbf{x}V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)\,.$ (27)
Note that since every intersection is counted only once, compared with (24) the $1/2$ factor is gone and the domain of integration becomes $\mathcal{B}$ instead of $\partial\mathcal{M}$.
@@
We see that, only the threads in the third class intersect with $\mathcal{E}_{\mathcal{A}}$ once. In other words, for a specific area element $\mathrm{d}\Sigma$ embedded in $\mathcal{E}_{\mathcal{A}}$, all of the PEE threads passing through $\mathrm{d}\Sigma$ from the inside to the outside are given by the PEE threads emanating from $\mathcal{A}$ with the initial orientation pointing from the boundary to the bulk. This means the spherical region $\mathcal{A}$ determined by $\mathrm{d}\Sigma$ is exactly the region $\mathcal{B}$ which satisfies (26). See Fig. 4 for an explicit illustration. Then using the theorem 1 (or (13)) for spherical regions, we arrive at
@@
which eventually leads to the proof of the theorem 2 and our reformulation (21) of the RT formula.
One may wonder whether the above derivation still is valid, if we embed $\mathrm{d}\Sigma$ on the RT surface of a non-spherical region, for example a strip. In Fig. 5 we show the PEE threads passing through the RT surface for a strip region. We see that there are PEE threads emanating from the strip and passing through $\mathcal{E}_{A}$ twice, one from the inside to the outside and the other from the outside to the inside. For any $\mathrm{d}\Sigma$ on the RT surface of the strip, the whole set of PEE threads emanating from the strip with the initial orientation pointing inward the bulk does not satisfy (26). That is, the set of PEE threads emanating from the strip does not all pass through $d\Sigma$ from one side to the other, hence
density of $\textit{density of intersections}~{}~{}\neq~{}~{}\int_{\rm strip}\mathrm{d}^{d-1}\textbf{x}V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)\,,$ (29)
@@
Figure 5: The RT surface $\mathcal{E}_{A}$ for a strip region $A$ in Poincaré AdS4. The PEE threads (the gray curves) with both endpoints inside or outside $A$ can possibly pass through $\mathcal{E}_{A}$ twice.
@@ 4 Reconstruction beyond the RT surfaces
Previously we proved the theorem 2, and use it to give a complete reformation of the RT formula in AdS space. Actually, the theorem 2 is very powerful and enough to generalize our reconstruction from the RT surfaces to the area of a generic co-dimension 2 hyper-surface in the AdS bulk. [^4] By co-dimension 2 hypersurfaces we mean co-dimension 2 hypersurfaces in AdS spacetime. In other words, we can interpret the area of all the geometric quantities in terms of the boundary PEEs. The generalization is straightforward, given an arbitrary co-dimension 2 surface $\Sigma$ we have
$\frac{\mathrm{Area}[\Sigma]}{4G}=\mathcal{N}(\Sigma)=\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{y}~{}\omega_{\Sigma}(\textbf{x},\textbf{y})\mathcal{I}(\textbf{x},\textbf{y})\,.$ (30)
@@
[…]
@@ 5 Equivalence to the Crofton formula
The main result (30) in this paper indeed indicates a pure mathematical theorem if we cancel the $1/(4G)$ coefficient [^5] Note that there is a factor $c/6=1/(4G)$ in $\mathcal{I}(\textbf{x},\textbf{y})$. on both sides. It claims that, when the density of geodesics in AdS space is properly defined, one can calculate the area of any co-dimension 2 surface by counting the number of intersections between the surface and all the geodesics in AdS. Remarkably, similar mathematical statements for a generic Riemannian manifold has been developed in integral geometry (see [67] for systematic introduction), which is called the Crofton formula. In this section we will first review the Crofton formula, then we show our result (30) is equivalent to the Crofton formula.
5.1 The Crofton formula
where $\overline{\text{Gr}}_{1}(\mathbb{R}^{2})$ is the space of geodesics in $\mathbb{R}^{2}$, $\mathrm{d}\Gamma$ is an appropriately normalized measure invariant under rigid motions, and #$\#(C\cap\Gamma)$ is the number of times a geodesic $\Gamma$ intersects with the curve $C$. The geodesics in $\mathbb{R}^{2}$ can be parameterized by
$x\cos\phi+y\sin\phi-p=0,$ (33)
$\mathrm{d}\Gamma=\mathrm{d}p\wedge\mathrm{d}\phi.$ (34)
It is easy to check this measure is invariant under rigid motions. Then the Crofton formula can be written as  [^6] For oriented straight lines, the range of $p$ is (,)($\infty$,$\infty$)( - ∞ , ∞ ) and the factor in front of the integral is $\frac{1}{4}$.
$\text{Length}(C)=\frac{1}{2}\int^{\infty}_{0}\mathrm{d}p\int^{2\pi}_{0}\mathrm{d}\phi\#(C\cap\Gamma(p,\phi)).$ (35)
Assuming that $C$ is parameterized by $x=x(s),y=y(s)$, where the parameter $s$ is the length parameter. The number of intersections can be counted along the curve. Let $\alpha$ denote the angle between the tangent at the point on $C$ and the straight line $\Gamma$ intersect $C$ at this point and $\theta$ the angle between the tangent to $C$ and $x$-axis. We have
$\phi=\alpha+\theta-\frac{\pi}{2}.$ (36)
Now we use $s$ and $\alpha$ to express $\mathrm{d}\Gamma$. By (33) and (36), we have
$\mathrm{d}p$ $=\cos\phi\mathrm{d}x+\sin\phi\mathrm{d}\phi+(-x\sin\phi+y\cos\phi)\mathrm{d}\phi,$ (37)
$\mathrm{d}\phi$ $=\mathrm{d}\alpha+\frac{\mathrm{d}\theta}{\mathrm{d}s}\mathrm{d}s,$
and because $\mathrm{d}x=\cos\theta\mathrm{d}s,\mathrm{d}y=\sin\theta\mathrm{d}s$, we get
$\mathrm{d}\Gamma=|\sin\alpha|\mathrm{d}s\wedge\mathrm{d}\alpha.$ (38)
The right-hand side of (35) is then
$\phi$$p$$\Gamma$$y$$x$$C$$\theta$$\alpha$
@@ 5.1 The Crofton formula
The Crofton formula can be generalized to Riemannian manifolds. Let us consider a $d$-dimensional Riemannian manifold $M$ (with a metric $g_{\mu\nu}$) and an arbitrary $(d-1)$-dimensional immersed hypersurface $\Sigma\subset M$, then
where $\text{Geod}(M)$ is the space of all geodesics in $M$, and $\mathrm{d}\Gamma$ is a properly chosen measure on $\text{Geod}(M)$ and $c$ is a constant. The measure $\mathrm{d}\Gamma$, following [68], is $|\omega|^{n-1}$ where $\omega$ is the symplectic form on $\text{Geod}(M)$ inherited from the canonical symplectic form $\omega_{M}$ on cotangent bundle $T^{*}M$ of $M$ in the following way:
${S^{*}M}$${T^{*}M}$${\text{Geod}(M)}$$\scriptstyle{i}$$\scriptstyle{\pi}$
where $S^{*}M$ denotes the unit co-sphere bundle of $M$; the canonical projection $\pi$ sends a given unit co-vector to the geodesic with this co-vector as initial condition and $i:S^{*}M\rightarrow T^{*}M$ is the canonical inclusion. Then there is a unique symplectic form $\omega$ on $\text{Geod}(M)$ such that $\pi^{*}\omega=i^{*}\omega_{M}$. The basic ideas behind the Crofton formula (40) are the co-area formula and symplectic reduction. Following [67], we present an intuitive derivation of the Crofton formula using the similar strategy as in the $\mathbb{R}^{2}$ case. We decompose the measure $\mathrm{d}\Gamma$ into two parts, one of which represents the position of all the points on the hypersurface $\Sigma$ and the other represents all the tangent direction of the geodesics passing through any point on $\Sigma$.
$\mathcal{L}=\sqrt{g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}},$ (41)
The co-sphere bundle $S^{*}M$ (i.e. pair $(x^{\mu},p_{\mu})$ with $p^{2}=1$) forms a manifold with $\dim=2d-1$, which is the phase space of $M$. And $(x^{\mu},p_{\mu})$ serves as a local coordinate system for $S^{*}M$, called the canonical coordinates. The geodesics $\Gamma$ of ${M}$ define a foliation $F_{\Gamma}$ of $S^{*}M$, in which the leaves $S^{*}M/F_{\Gamma}$ have dimension $22d-$ 2. The canonical two-form
$\mathrm{d}p_{\mu}\wedge\mathrm{d}x^{\mu}$ (43)
on $S^{*}M$ is invariant under local coordinate transformations and the displacements on the leaves of the foliation (i.e. independent of the parameter $\tau$) [67]. It follows that the following $(2d-2)$-form on $S^{*}M$ is also invariant
$\mathrm{d}\Gamma$ $=(\mathrm{d}p_{\mu}\wedge\mathrm{d}x^{\mu})^{d-1}$ (44)
which defines an invariant measure of $(2d-2)$-dimensional set of geodesics $\Gamma$. It is called kinematic measure. Also the space of all the geodesics in the manifold is called the kinematic space [21, 22, 23, 69].
In other words, the bulk geodesics are parameterized by the position where they intersect with $\Sigma$ and the intersection direction captured by the $d-1$ independent conjugate momentum. To perform the integral over $\mathrm{d}\Gamma$, note that the conjugate momentum can be parameterized by the angle $\alpha_{i}$ between the momentum and $x^{i}$ axis: $p_{i}=\sqrt{g_{ii}}\cos\alpha_{i}$. Up to the sign, the kinematic measure becomes
where we recognize the area element $\mathrm{d}\sigma$ of a $(d-1)$-dimensional hypersurface $\Sigma$:
and the area element on the unit $(d-1)$-dimensional sphere in the direction $p_{d}$
Hence we can write
and perform integral for the form $\mathrm{d}\Gamma$ over $\Sigma$ and the upper half of the $(d-1)$-dimensional unit sphere777 The integral is over half the unit sphere because the geodesics are unoriented.
@@
$=\mathrm{Area}(\Sigma)\frac{\Omega_{d-2}}{d-1},$
where $\Omega_{d-2}=\frac{2\pi^{(d-1)}/2}{\Gamma((d-1)/2)}$ is the area of the $(d-2)$-dimensional sphere and in the last step we apply the following relation
5.2 The kinematic measure from the PEE structure
@@
The key feature of pure AdS is that all geodesics have two endpoints on the boundary, and there is one and only one geodesic connecting a given pair of boundary points. Therefore, we are free to parameterize the geodesics using its endpoints on the boundary. Now we restrict $x_{1}$ and $x_{2}$ to the boundary, and parameterize the geodesics using its endpoints $(\textbf{x}_{1},\textbf{x}_{2})$, with coordinates $x_{1,2}^{i},i=1,\ldots,d-1$. The invariant density of geodesics is given by the $(2d-2)$-form (44). In terms of boundary coordinates $(\textbf{x}_{1},\textbf{x}_{2})$, the density becomes
@@
which is exactly (30). Here $\omega_{\Sigma}(\textbf{x}_{1},\textbf{x}_{2})$ is just #$\#(\Sigma\cap\Gamma)$ with $\Gamma$ being the PEE thread connecting the pair of boundary points $\{\textbf{x}_{1},\textbf{x}_{2}\}$. Hence our main result (30) is equivalent to the Croftron formula in AdS space.
@@
$\frac{\mathrm{Area}(\Sigma)}{4G}=\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}_{1}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}_{2}\omega_{\Sigma}(\textbf{x}_{1},\textbf{x}_{2})\det\left(\frac{\partial^{2}\ell(\textbf{x}_{1},\textbf{x}_{2})}{\partial\textbf{x}_{1}\partial\textbf{x}_{2}}\right).$ (59)

### v4

@@ 3 Reformulation of the RT formula
$\boxed{\mathcal{N}(\Sigma_{A})=\frac{1}{2}\int_{\Sigma_{A}}\mathrm{d}\Sigma_{A}\sqrt{h}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}|V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)|\,.}$ (23)
Here $\mathrm{d}\Sigma_{A}\sqrt{h}$ is the area of an infinitesimal area element on $\Sigma_{A}$, and the coefficient $1/2$ appears because we integrate x over the whole boundary such that the PEE thread connecting any two boundary point is doubly counted. The equivalence between (22) and (23) is guaranteed by the divergenceless property of the vector fields $V_{\textbf{x}}^{\mu}$  [^4] Usually it will be formidable to compute (22) for a generic $\Sigma_{A}$, as it will be hard to trace the weight for all of the threads. Here we give a simple example where both of (23) and (22) can be computed, and the results conincide. Let us consider the two-interval case with a disconnected entanglement wedge shown in Fig.3. Given the RT surface $\mathcal{E}_{A}=\mathcal{E}_{A_{1}}\cup\mathcal{E}_{A_{2}}$, it is easy to count the weight for each thread, and the computation of (22) is exactly given by (17), which gives $\mathcal{N}(\mathcal{E}_{A})=S_{A_{1}}+S_{A_{2}}=\frac{\mathrm{Area}[\mathcal{E}_{A_{1}}]+\mathrm{Area}[\mathcal{E}_{A_{2}}]}{4G}$, where we have used the normalization property for single intervals. On the other hand, Since both of $\mathcal{E}_{A_{1}}$ and $\mathcal{E}_{A_{2}}$ are RT surfaces for single intervals, the statement 1 applies hence the density of intersections between these two surfaces and the PEE network is given by the constant $\frac{1}{4G}$. Then we also get $\mathcal{N}(\mathcal{E}_{A})=\mathcal{N}(\mathcal{E}_{A_{1}}\cup\mathcal{E}_{A_{2}})=\frac{\mathrm{Area}[\mathcal{E}_{A_{1}}]+\mathrm{Area}[\mathcal{E}_{A_{2}}]}{4G}$. .
According to (23), the density of intersections between $\Sigma_{A}$ and the PEE network at any point $Q$ on $\Sigma_{A}$ is just given by
$\textit{density of intersections}~~=~~\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}|V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)|\,.$ (24)
Since $\Sigma_{A}$ is an arbitrary homologous surface, given any bulk point $Q$ and any normal direction $n_{\mu}(Q)$, if we can calculating the above integration and prove
$\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}|V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)|=\frac{1}{4G}\,$, (25)
then the statement 2 follows. For example, in Poincaré AdS3 as Fig. 4, (24) sums over all the PEE threads that pass through $\Sigma_{A}$ at $\mathrm{d}\Sigma|_{Q}$, shown as the gray geodesics in Fig. 4. The factor $(1/2)$ accounts for the double counting of each geodesic. And (25) is to say that such a density is a universal constant $1/4G$. Detailed explaination and proof of (25) go as follows.
Let us consider an arbitrary area element $\mathrm{d}\Sigma$ at any point $Q$ in the bulk and the associated normal unit vector $n^{\mu}(Q)$, directly calculating the integration (24) is complicated, because of the absolute symbol in the integrand. We always start from certain configuration of orientations for all the threads as we use vector fields to describe them in (24). Nevertheless, we stress that the PEE threads are un-oriented lines, and the orientation come with the vector field description are redundant information, which are excluded by the absolute value sign. For an arbitrary configuration of the orientations, it is also formidable to compute (24) due to the absolute value sign. Given an infinitesimal area element in the bulk, the trick of the computation is to find a proper assignment for the orientations of the threads passing through it, such that all the threads point from one same side of the area element to the other side, hence the cancellation due to opposite orientations can be avoided. This allows us to drop the absolute value sign and the number of intersections equals to the flux passing through the area element.
More explicitly, our strategy is to find the boundary region $\mathcal{B}$ such that, for any $\textbf{x}\in\mathcal{B}$ we have
@@
$\textit{density of intersections}~~=~~\int_{\mathcal{B}}\mathrm{d}^{d-1}\textbf{x}V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)\,.$ (27)
Note that since every intersection is counted only once, compared with (24) the $1/2$ factor is gone and the domain of integration becomes $\mathcal{B}$ instead of $\partial\mathcal{M}$. Although the analysis for static spherical regions [66] looks quite special, it contains the key ingredient to prove (25).
@@
We see that only the threads in the third class intersect with $\mathcal{E}_{\mathcal{A}}$ once. In other words, for a specific area element $\mathrm{d}\Sigma$ embedded in $\mathcal{E}_{\mathcal{A}}$, all of the PEE threads passing through $\mathrm{d}\Sigma$ from the inside to the outside are given by the PEE threads emanating from $\mathcal{A}$ with the initial orientation pointing from the boundary to the bulk. This means the spherical region $\mathcal{A}$ determined by $\mathrm{d}\Sigma$ is exactly the region $\mathcal{B}$ which satisfies (26). See Fig. 4 for an explicit illustration. Then using the statement 1 (or (13)) for spherical regions, we arrive at
@@
which eventually leads to the proof of the statement 2 and our reformulation (21) of the RT formula. We emphasize that the embeding of $\mathrm{d}\Sigma$ into $\mathcal{E}_{\mathcal{A}}$ is purely out of computational convenience, which allows us to remove the absolute sign in the integration and perform it with ease. In principle, (25) is independent of the assignment of the orientations of $V_{\textbf{x}}^{\mu}$.
One may wonder whether the above derivation still is valid, if we embed $\mathrm{d}\Sigma$ on the RT surface of a non-spherical region $\cal A$ and assign positive direction to these threads from $\mathcal{A}$ to $\bar{\cal A}$, for example in a strip. The answer is no. In Fig. 5 we show the PEE threads passing through the RT surface for a strip region. We see that there are PEE threads emanating from the strip and passing through $\mathcal{E}_{A}$ twice, one from the inside to the outside and the other from the outside to the inside. For any $\mathrm{d}\Sigma$ on the RT surface of the strip, the whole set of PEE threads emanating from the strip with the initial orientation pointing inward the bulk does not satisfy (26). That is, the set of PEE threads emanating from the strip does not all pass through $d\Sigma$ from one side to the other, hence we cannot remove the absolute value in the integration:
$\textit{density of intersections}=\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}|V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)|\,~~\neq~~\int_{\rm strip}\mathrm{d}^{d-1}\textbf{x}V_{\textbf{x}}^{\mu}(Q)n_{\mu}(Q)\,$, (29)
@@
Figure 5: The RT surface $\mathcal{E}_{A}$ for a strip region $A$ in Poincaré AdS4. The PEE threads (the gray curves) with both endpoints inside or outside AA can possibly pass through $\mathcal{E}_{A}$ twice.
@@ 4 Reconstruction beyond the RT surfaces
Previously we proved the statement 2, and use it to give a reformation of the RT formula in AdS space. Actually, the statement 2 is very powerful and enough to generalize our reconstruction from the RT surfaces to the area of a generic co-dimension 2 hyper-surface in the AdS bulk. [^5] By co-dimension 2 hypersurfaces we mean co-dimension 2 hypersurfaces in AdS spacetime. In other words, we can interpret the area of all the geometric quantities in terms of the boundary PEEs. The generalization is straightforward, given an arbitrary co-dimension 2 surface $\Sigma$ we have
$\frac{\mathrm{Area}[\Sigma]}{4G}=$ $\mathcal{N}(\Sigma)$
$=$ $\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{y}~\omega_{\Sigma}(\textbf{x},\textbf{y})\mathcal{I}(\textbf{x},\textbf{y})\,.$ (30)
@@
Interestingly, given the position of an infinitesimal area element $\mathrm{d}\Sigma$ at $Q$, all the PEE threads passing through $Q$ contribute to the reconstruction of $\mathrm{d}\Sigma$. It is easy to see that, the scale $|\textbf{x}-\textbf{y}|$ of all the two-point PEEs (or PEE threads) $\mathcal{I}(\textbf{x},\textbf{y})$ that contribute to the reconstruction of $\mathrm{d}\Sigma$, is lower bounded. For example, in Poincaré AdSd+1, the scale of those two-point PEEs satisfies
@@ 5 Equivalence to the Crofton formula
The main result (4) in this paper indeed indicates a pure mathematical theorem if we cancel the $1/(4G)$ coefficient [^6] Note that there is a factor $c/6=1/(4G)$ in $\mathcal{I}(\textbf{x},\textbf{y})$. on both sides. It claims that, when the density of geodesics in AdS space is properly defined, one can calculate the area of any co-dimension 2 surface by counting the number of intersections between the surface and all the geodesics in AdS. Remarkably, similar mathematical statements for a generic Riemannian manifold have been developed in integral geometry, which is called the Crofton formula  [^7] See A for more details of the Crofton formula. A comprehensive review of the classical integral geometry is provided in [81].
@@
The Crofton formula states that given a $d$-dimensional Riemannian manifold $M$ and a (d1)(d-1)-dimensional hypersurface $\Sigma$, then the area of $\Sigma$ is given by
@@
$\frac{1}{2}\frac{d-1}{\Omega_{d-2}}\int_{\text{Geod}(M)}\#(\Sigma\cap\Gamma)\mathrm{d}\Gamma=\mathrm{Area}(\Sigma),$ (32)
where $\Omega_{d-1}=\frac{2\pi^{{d-1}/2}}{\Gamma((d-1)/2)}$ is the area if the (d2)(d-2)-dimensional sphere, $(\Sigma\cap\Gamma)$ is the number of intersections of $\Sigma$ and a geodesic $\Gamma$ and $\mathrm{d}\Gamma$ is a properly chosen measure over the space of geodesics $\text{Geod}(M)$. The measure $\mathrm{d}\Gamma$ has to be invariant under changes of local coordinates and invariant under translations along geodesics. More explicitly, given the metric $g_{\mu\nu}$ of the Riemannian manifold $M$, we put
$\mathcal{L}=\sqrt{g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}},\quad p_{\mu}=\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}},$ (33)
and $\dot{x}^{\mu}:=\mathrm{d}x^{\mu}(\tau)/\mathrm{d}\tau$, with some parameters $\tau$ of geodesics. Then the measure $d\Gamma$ is
$\mathrm{d}\Gamma=(\mathrm{d}p_{\mu}\wedge\mathrm{d}x^{\mu})^{d-1}.$ (34)
Now we will show our result (4) is equivalent to the Crofton formula.
@@
The key feature of pure AdS is that all geodesics have two endpoints on the boundary, and there is one and only one geodesic connecting a given pair of boundary points. Therefore, we are free to parameterize the geodesics using its endpoints on the boundary. Now we restrict $x_{1}$ and $x_{2}$ to the boundary, and parameterize the geodesics using its endpoints $(\textbf{x}_{1},\textbf{x}_{2})$, with coordinates $x_{1,2}^{i},i=1,\ldots,d-1$. The invariant density of geodesics is given by the (2d2)(2d-2)-form (51). In terms of boundary coordinates $(\textbf{x}_{1},\textbf{x}_{2})$, the density becomes
@@
which is exactly (4). Here $\omega_{\Sigma}(\textbf{x}_{1},\textbf{x}_{2})$ is just $(\Sigma\cap\Gamma)$ with $\Gamma$ being the PEE thread connecting the pair of boundary points $\{\textbf{x}_{1},\textbf{x}_{2}\}$. Hence our main result (4) is equivalent to the Croftron formula in AdS space.
@@
$\frac{\mathrm{Area}(\Sigma)}{4G}=\frac{1}{2}\frac{d-1}{\Omega_{d-2}}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}_{1}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}_{2}\omega_{\Sigma}(\textbf{x}_{1},\textbf{x}_{2})\det\left(\frac{\partial^{2}\ell(\textbf{x}_{1},\textbf{x}_{2})}{\partial\textbf{x}_{1}\partial\textbf{x}_{2}}\right).$ (40)